OpenStudy (anonymous):

I need help with these 4 math problems please tell someone can help me answer them -2x-2y=10 for y A=l w for W A=1/2bh for b A=1/2 (a+b)h for a

OpenStudy (anonymous):

arrange the equation according to the variable they asked you..

OpenStudy (anonymous):

y=-x-5 w=A/l b=2A/h a=(2A/h)-b

OpenStudy (anonymous):

It's just rearrangement of variables...if you need work/help understanding let me know & I'll show you!